### v1

# GUI Agents: A Survey

Abstract
Graphical User Interface (GUI) agents, powered by Large Foundation Models, have emerged as a transformative approach to automating human-computer interaction. These agents autonomously interact with digital systems or software applications via GUIs, emulating human actions such as clicking, typing, and navigating visual elements across diverse platforms. Motivated by the growing interest and fundamental importance of GUI agents, we provide a comprehensive survey that categorizes their benchmarks, evaluation metrics, architectures, and training methods. We propose a unified framework that delineates their perception, reasoning, planning, and acting capabilities. Furthermore, we identify important open challenges and discuss key future directions. Finally, this work serves as a basis for practitioners and researchers to gain an intuitive understanding of current progress, techniques, benchmarks, and critical open problems that remain to be addressed.
1 Introduction
Large Foundation Models (LFMs) have significantly transformed both the landscape of AI research and day-to-day life (Bommasani et al., 2022; Kapoor et al., 2024; Schneider et al., 2024; Naveed et al., 2023; Wang et al., 2024d). Recently, we have witnessed a paradigm shift from using LFMs purely as conversational chatbots (Touvron et al., 2023; Chiang et al., 2023; Dam et al., 2024) to employing them for performing actions and automating useful tasks (Wang et al., 2024b; Zhao et al., 2023; Yao et al., 2023; Shinn et al., 2023; Shen et al., 2024b; Cheng et al., 2024c). In this direction, one approach stands out: leveraging LFMs to interact with digital systems, such as desktops and mobile phones, or software applications such as a web browser, through Graphical User Interfaces (GUIs) in the same way humans do—for example, by controlling the mouse and keyboard to interact with visual elements displayed on a device’s monitor Iong et al. (2024); Hong et al. (2023); Lu et al. (2024); Shen et al. (2024a).
This approach holds great potential, as GUIs are ubiquitous across almost all computer devices that humans interact with in their work and daily lives. However, deploying LFMs in such environments poses unique challenges, such as dynamic layouts, diverse graphical designs across different platforms, and grounding issues—for instance, fine-grained recognition of elements within a page that are often small, numerous, and scattered Liu et al. (2024b). Despite these challenges, many early efforts have shown significant promise (Lin et al., 2024; Cheng et al., 2024a), and growing interest from major players in the field is becoming evident [^1] Anthropic, Google DeepMind, OpenAI.
Given the immense potential and rapid progress in this field, we propose a unified and systematic framework to categorize the various types of contributions within this space.
Organization of this Survey. We begin our survey by clearly defining the term “GUI Agent,” followed by a traditional RL formalism of GUI Agent tasks in Section 2. We then summarize different datasets and environments in Section 3 to provide readers a clearer picture of the kinds of problem settings currently available. We summarize various GUI Agent architectural designs in Section 4, followed by different ways of training them in Section 5. Lastly, we discuss open problems and future prospects of GUI Agent research in Section 6.
2 Preliminaries
Definition 1 (GUI Agent).
An intelligent autonomous agent that interacts with digital platforms, such as desktops, or mobile phones, through their Graphical User Interface. It identifies and observes interactable visual elements displayed on the device’s screen and engages with them by clicking, typing, or tapping, mimicking the interaction patterns of a human user.
Problem Formulation.
GUI Agent involves an agent interacting with an environment in a sequential manner. The environment can generally be modeled as a Partially Observable Markov Decision Process (POMDP), defined by a tuple $(\mathcal{U},\mathcal{A},\mathcal{S},\mathcal{O},T)$, where $\mathcal{U}$ is the task space, $\mathcal{A}$ is the action space, $\mathcal{S}$ is the state space, $\mathcal{O}$ is the observation space, and $T:\mathcal{S}\times\mathcal{A}\to\mathcal{P}(\mathcal{S})$ is a state transition function mapping a state-action pair to a probability distribution over subsequent states.
A GUI Agent is a mapping $\pi:\mathcal{O}\to\mathcal{A}$. Given a task $u\in\mathcal{U}$, the agent proceeds through a sequence of actions to complete the task. At each step $t$, based on the available observation $o\in\mathcal{O}$, the agent $\pi$ must predict the next action $a\in\mathcal{A}$. The environment then transitions to a new state $s^{\prime}\in\mathcal{S}$ according to $T$. Depending on the design of the environment, the agent may receive a reward $r=R(s,a,s^{\prime})$, where $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a reward function.
3 Benchmarks
GUI agents are developed and evaluated on various platforms, including desktops, mobile phones, and web browser environments. This section summarizes benchmarks for all of these platform types.
When evaluating GUI Agents, it is crucial to distinguish between an environment and a dataset. A dataset is a static collection of data point, where each consists of several input features (e.g., a question, a screenshot of the environment, or the current state of the environment) and some output features (e.g., correct answers or actions to be taken). A dataset remains unchanged throughout the evaluation process. In contrast, an environment is an interactive simulation that represents a real-world scenario of interest. A GUI environment includes the GUI interface of a mobile phone or a desktop. Unlike datasets, environments are dynamic—actions taken within the environment can alter its state, hence, allowing modeling the problem as Markov Decision Processes (MDPs) or Partially Observable MDPs (POMDPs), with defined action, state, and observation spaces, and a state transition function.
Another critical dimension of the existing benchmarks for GUI Agentsis the distinction between the open-world and closed-world assumptions. Closed-world datasets or environments presume that all necessary knowledge for solving a task is contained within the benchmark itself. In contrast, open-world benchmarks relax this constraint, allowing relevant information required to complete a task to exist outside the benchmark.
3.1 Static Datasets
3.1.1 Closed-World Datasets
RUSS dataset introduces real-world instructions mapped to a domain-specific language (DSL) that enables agents to execute web-based tasks with high precision (Xu et al., 2021). Similarly, Mind2Web expands the task set to 2000 diverse tasks (Deng et al., 2023), and MT-Mind2Web adapts into conversational settings with multi-turn interactions (Deng et al., 2024). In contrast, TURKINGBENCH focuses on common micro tasks in crowdsourcing platforms, featuring a rich mix of textual instructions, multi-modal elements, and complex layouts (Xu et al., 2024). Focusing on visual and textual interplay, VisualWebBench includes OCR, element grounding, and action prediction tasks, which require fine-grained multimodal understanding (Liu et al., 2024b). Similarly, ScreenSpot focuses on GUI grounding for clicking and typing directly from screenshots (Cheng et al., 2024b). Complementing this, WONDERBREAD extends evaluation to business process management tasks, emphasizing workflow documentation and improvement rather than automation alone (Wornow et al., 2024). EnvDistraction dataset explores agent susceptibility to distractions in GUI environments, offering insights into faithfulness and resilience under cluttered and misleading contexts (Ma et al., 2024). NaviQAte introduces functionality-guided web application navigation, where tasks are framed as QA problems, pushing agents to extract actionable elements from multi-modal inputs (Shahbandeh et al., 2024).
Evaluating on static closed-world datasets is particularly convenient, thanks to their lightweight and ease in setting up compared to environments. They are also especially valuable for fine-grained evaluation, reproducibility, and comparing models under identical conditions. However, they lack the dynamism of real-world applications, as models are tested on fixed data rather than adapting to new inputs or changing scenarios.
3.1.2 Open-World Datasets.
While most existing datasets are designed under the closed-world assumption, several datasets do not follow this paradigm. GAIA dataset tests agent integration diverse modalities and tools to answer real-world questions, often requiring web browsing or interaction with external APIs (Mialon et al., 2023). WebLINX emphasizes multi-turn dialogue for interactive web navigation on real-world sites, enhancing agents’ adaptability and conversational skills (Lù et al., 2024).
Evaluation on static open-world datasets balances the ease of setting up an evaluation setting with realism since the agents interact with real-world websites. However, due to the nature of real-world websites, they are often unpredictable and prone to changes, which makes it more challenging to reproduce and compare with prior methods.
3.2 Interactive Environments
3.2.1 Closed-World Environments.
Closed-world interactive environments provide controlled and reproducible settings for evaluating agent capabilities. MiniWoB offers synthetic web tasks requiring interactions with webpages using mouse and keyboard inputs (Shi et al., 2017). It focuses on fundamental skills like button clicking and form filling, providing a baseline for evaluating low-level interaction. CompWoB extends MiniWoB with compositional tasks, requiring agents to handle multi-step workflows and generalize across task sequences (Furuta et al., 2023). This introduces dynamic dependencies that reflect real-world complexity. WebShop simulates e-shopping tasks that challenge agents to navigate websites, process instructions, and make strategic decisions (Yao et al., 2022). WebArena advances realism with self-hosted environments across domains like e-commerce and collaborative tools, requiring agents to manage long-horizon tasks (Zhou et al., 2023b). VisualWebArena adds multimodal challenges, integrating visual and textual inputs for tasks like navigation and object recognition (Koh et al., 2024a). Shifting to enterprise settings, WorkArena evaluates agent performance in complex UI environments, focusing on knowledge work tasks in ServiceNow platform (Drouin et al., 2024). ST-WebAgentBench incorporates safety and trustworthiness metrics, assessing policy adherence and minimizing risky actions, critical for business deployment (Levy et al., 2024). Lastly, VideoWebArena introduces long-context video-based tasks, requiring agents to understand instructional videos and integrate them with textual and visual data to complete tasks. It emphasizes memory retention and multimodal reasoning (Jang et al., 2024).
Closed-world environments serve as evaluation platforms that mimic the dynamism of real-world environments while offering stability and reproducibility. However, setting up such benchmarks is often challenging, as they typically require considerable storage space and engineering skills.
3.2.2 Open-World Environments.
Open-world interactive environments challenge agents to navigate dynamic, real-world websites with evolving content and interfaces. WebVLN introduces a novel benchmark for vision-and-language navigation on websites, requiring agents to interpret visual and textual instructions to complete tasks such as answering user queries (Chen et al., 2024). It emphasizes multimodal reasoning by integrating HTML structure with rendered webpages, setting a foundation for realistic web navigation. WebVoyager leverages LLM to perform end-to-end navigation on 15 real websites with diverse tasks (He et al., 2024b). Its multimodal approach integrates screenshots and HTML content, enabling robust decision-making in dynamic online settings. AutoWebGLM optimizes web navigation through HTML simplification and reinforcement learning (Lai et al., 2024). This framework tackles the challenges of diverse action spaces and complex web structures, demonstrating significant improvement in real-world tasks with its AutoWebBench benchmark. MMInA evaluates agents on multihop, multimodal tasks across evolving real-world websites (Zhang et al., 2024e). The benchmark includes 1,050 tasks requiring sequential reasoning and multimodal integration to complete compositional objectives, such as comparing products across platforms. WebCanvas pioneers a dynamic evaluation framework to assess agents in live web environments (Pan et al., 2024). Its Mind2Web-Live dataset captures the adaptability of agents to interface changes and includes metrics like key-node-based intermediate evaluation, fostering progress in online web agent research.
Open-world environments are ideal for achieving both realism and dynamism. However, getting consistent evaluation and reproducibility is difficult as they evaluate agents on live websites that are subject to frequent changes.
3.3 Evaluation Metrics
Task Completion Metrics.
The majority of benchmarks use task completion rate as the primary metric to measure GUI Agents’ performance. However, different papers define task completion differently. Success can be defined as whether an agent successfully stops at a goal state (Chen et al., 2024; Zhou et al., 2023b), with Zhou et al. (2023b) programmatically checking if the intended outcome has been achieved (e.g., a comment has been posted, or a form has been completed), or whether the returned results exactly match the ground truth labels (Shi et al., 2017; Yao et al., 2022; Koh et al., 2024a; Drouin et al., 2024; Levy et al., 2024; Mialon et al., 2023). Another approach is to measure success based on whether an agent completes all required subtasks (Lai et al., 2024; Zhang et al., 2024e; Pan et al., 2024; Furuta et al., 2023; Jang et al., 2024; Cheng et al., 2024b). This approach can be further extended to measure partial success, as shown in Zhang et al. (2024e). WebVoyager uses GPT-4V to automatically determine success based on the agent’s trajectory, reporting a high agreement rate of 85.3% with human judgments (He et al., 2024b). Instead of using a single final-state success metric, WebLINX measures an overall success rate based on aggregated turn-level success rates across tasks (Lù et al., 2024). The turn-level success rates are computed depending on the type of actions, e.g., Intersection Over Union (IoU) for click or submit actions, and F1 for say or textinput actions. Lastly, there are task-specific metrics to measure success, e.g., using ROUGE-L, F1 for open-ended generation (Liu et al., 2024b; Xu et al., 2024; Wornow et al., 2024), accuracy for multiple choice question tasks (Liu et al., 2024b), Precision and Recall for Standard Operating Procedure (SOP) validation (Wornow et al., 2024), and so on.
Intermediate Step Metrics.
While the task completion rate is a straightforward single-numeric metric that simplifies comparing the overall performance of agents, it fails to provide clear insights into their specific behaviors. Although some fine-grained metrics measure step-wise performance, their scope remains limited. WebCanvas evaluates step scores using three distinct targets: URL Matching, which verifies whether the agent navigated to the correct webpage; Element Path Matching, which checks if the agent interacted with the appropriate UI element, such as a button or text box; and Element Value Matching, which ensures the agent inputted or extracted the correct values, such as filling a form or reading text. WebLINX uses an intent match metric to assess whether the predicted action’s intent aligns with the reference intent. Similarly, Mind2Web and MT-Mind2Web evaluate Element Accuracy by measuring the rate at which the agent selects the correct elements. These systems also measure the precision, recall, and F1 score for token-level operations, such as clicking or typing, and calculate the Step Success Rate, which reflects the proportion of individual task steps completed correctly. While step-wise evaluations provide more fine-grained insight into the agent’s performance, it is often challenging to collect reference labels at the step level while also providing enough flexibility to consider different paths to achieve the original tasks.
Efficiency, Generalization, Safety and Robustness Metrics.
Lastly, we summarize additional metrics that evaluate various aspects of GUI agents beyond their raw performance. Existing benchmarks include metrics for efficiency (Shahbandeh et al., 2024; Chen et al., 2024; Shahbandeh et al., 2024), generalization across diverse or compositional task settings (Furuta et al., 2023), adherence to safety policies (Levy et al., 2024), and robustness to environmental distractions (Ma et al., 2024).
4 GUI Agent Architectures
This section focuses on various architectural designs of a GUI Agent agent, which we categorize into four main types: (1) Perception: designs that enable the GUI Agent agent to perceive and interpret observations from its environment; (2) Reasoning: designs related to the cognitive processes of a GUI Agent agent, such as using an external knowledge base for long-term memory access or a world model of the environment to support other modules like planning; (3) Planning: designs related to decomposing a task into subtasks and creating a plan for their execution; and (4) Acting: mechanisms that allow the GUI Agent agent to interact with the environment, including representing actions in natural language using specific templates, JSON, or programming languages as action representations.
4.1 Perception
Unlike API-based agents that process structured, program-readable data, GUI agents must perceive and understand the on-screen environment that is designed for human consumption. This requires carefully chosen interfaces that allow agents to discover the location, identity, and properties of the interactive elements. Broadly, these perception interfaces can be categorized into four types: accessibility-based, HTML/DOM-based, screen-visual-based, and hybrid ones, with each offering different capabilities and posing distinct privacy and implementation considerations.
4.1.1 Accessibility-Based Interfaces
Modern mobile and desktop operating systems usually provide accessibility APIs [^2] https://en.wikipedia.org/wiki/Computer˙accessibility that expose a semantic hierarchy of UI components, including their roles, labels, and states [^3] https://developer.apple.com/library/archive/documentation/Accessibility/Conceptual/AccessibilityMacOSX/OSXAXmodel.html [^4] https://developer.apple.com/design/human-interface-guidelines/accessibility [^5] https://learn.microsoft.com/en-us/windows/apps/design/accessibility/accessibility. GUI agents can utilize accessibility APIs to identify actionable elements and derive semantic cues without relying solely on pixel-based detection. These interfaces are resilient to minor layout changes or styling updates; however, their effectiveness depends on proper implementation by developers. Accessibility APIs may also be limited when dealing with highly dynamic elements (e.g., custom drawing canvases or gaming environments) and may not natively expose visual content. Although these APIs help reduce the complexity of visually parsing the screen, the agent may need additional perception methods for full functionality. On the positive side, accessibility-based interfaces typically require minimal sensitive user data, thereby reducing privacy concerns.
4.1.2 HTML/DOM-Based Interfaces
For web GUIs, agents frequently utilize the Document Object Model (DOM) to interpret the structural layout of a page. The DOM provides a hierarchical representation of elements, allowing agents to locate targets like buttons or input fields based on tags, attributes, or text content. However, raw HTML data or DOM tree usually has redundant and noisy structure. Various methods are proposed to handle this. Mind2Web (Deng et al., 2023) utilizes a fine-tuned small LM to rank the elements in a page before the final prediction of action with a large LM, and WebAgent (Gur et al., 2023) uses a specialized model HTML-T5 to generate task-specific HTML snippets. AutoWebGLM (Lai et al., 2024) designs an algorithm to simplify HTML content. While HTML/DOM-based interfaces provide rich structural data, they require careful preprocessing and, in some cases, additional heuristics or trained models to locate and interpret key UI components accurately.
4.1.3 Screen-visual-based Interfaces
With advances in computer vision and multimodal LLM, agents can utilize screen-visual information, like screenshots, to perceive on-screen environment. OmniParser (Lu et al., 2024) utilizes an existing multimodal LLM (e.g., GPT-4V) to parse a screenshot into a structured representation of the UI elements. However, screen-visual-based perception introduces privacy concerns since entire screenshots may contain sensitive information. Additionally, computational overhead increases as models must handle high-dimensional image inputs. Despite these challenges, such interfaces are crucial for agents operating in environments where high-quality accessibility interfaces and DOM information are unavailable, or environments where dynamic or visual information is crucial, like image or video editing software.
4.1.4 Hybrid Interfaces
To achieve robust and flexible performance across diverse environments, many GUI agents employ a hybrid approach. These systems combine accessibility APIs, DOM data, and screen-visual information to form a more comprehensive understanding of the interface. Leading methods in GUI agent tasks, such as OS-Atlas(Wu et al., 2024b) and UGround (Gou et al., 2024), demonstrates that hybrid interfaces that combine visual and textual inputs can enhance performance. Hybrid interfaces based approaches also facilitate error recovery—when accessibility or DOM data are incomplete or misleading, the agent can fall back on screen parsing, and vice versa.
4.2 Reasoning
WebPilot employs a dual optimization strategy for reasoning (Zhang et al., 2024d). WebOccam improves reasoning by refining the observation and action space of LLM agents (Yang et al., 2024). OSCAR introduces a general-purpose agent to generate Python code from human instructions (Wang and Liu, 2024). LAST leverages LLMs for reasoning, acting, and planning (Zhou et al., 2023a).
4.3 Planning
Planning involves decomposing a global task into multiple subtasks that progressively approach the goal state starting from an initial state Huang et al. (2024). Traditional planning methods, such as symbolic approaches and reinforcement learning, have significant limitations: symbolic methods require extensive human expertise to define rigid system rules and lack error tolerance Belta et al. (2007); Pallagani et al. (2022), while reinforcement learning demands impractical volumes of training data, often derived from costly environmental interactions Acharya et al. (2023). Recent advancements in LLM-powered agents offer a transformative alternative by positioning LLM-powered agents as the cognitive core for planning agents Huang et al. (2024). When equipping agents with GUIs as the medium, LLM-powered agents can directly interact with nearly all application domains and resources to enhance planning strategies. Based on what application domains/resources agents use for planning, we divide existing works into planning with internal and external knowledge.
4.3.1 Planning with Internal Knowledge
Planning with internal knowledge of GUI agents is to leverage the inherent knowledge to reason and think about the potential plans to fulfill the global task goals Schraagen et al. (2000). WebDreamer (Gu et al., 2024) uses LLMs to simulate the outcomes of the actions of each agent and then evaluate the result to determine the optimal plan at each step. MobA (Zhu et al., 2024) devises a two-level architecture to power the mobile phone management, with a high level for understanding user commands, tracking history memories and planning tasks, and a low level to act the planned module. Agent S (Agashe et al., 2024) introduces an experience-augmented hierarchical planning to perform complex computer tasks.
4.3.2 Planning with External Knowledge
Enabling LLM-powered agents to interact with diverse applications and resources through GUIs allows them to leverage external data sources, thereby enhancing their planning capabilities. For example, Search-Agent (Koh et al., 2024b) combines LLM inference with A* search to explore and backtrack to alternative paths explicitly, AgentQ (Putta et al., 2024) combines LLM with MCTS. Toolchain (Zhuang et al., 2023) models tool planning as a tree search algorithm and incorporates A* search to adaptively retrieve the most promising tool for subsequent use based on accumulated and anticipated costs. SGC (Wu et al., 2024a) decomposes the query and performs embedding similarity match between the concatenated subquery with the current retrieved task API and each of the existing APIs, and then selects the top one from the existing neighboring APIs. Thought Propagation Retrieval (Yu et al., 2023) prompts LLMs to propose a set of analogous problems and then applies established prompting techniques, like Chain-of-Thought, to derive solutions. The aggregation module subsequently consolidates solutions from these analogous problems, enhancing the problem-solving process for the original input. WebShop, Mind2Web, and WebArena Zhou et al. (2023c); Deng et al. (2023) allow agents to interact with webs to plan for web browsing for search. WMA (Chae et al., 2024) utilizes world models to address the mistakes made by LLMs for long-horizon tasks.
4.4 Acting
Acting in GUI agents involves translating the agent’s reasoning and planning outputs into executable steps within the GUI environment. Unlike purely text-based or API-driven agents, GUI agents must articulate their actions at a finer granularity—often down to pixel-level coordinates—while also handling higher-level semantic actions such as typing text, scrolling, or clicking on specific elements. Several directions of approaches have emerged:
Those utilizing textual interfaces may only rely on text-based metadata (HTML, accessibility trees) to identify UI elements. For example, WebAgent (Gur et al., 2023) and Mind2Web (Deng et al., 2023) use DOM or HTML representations to locate interactive elements. Similarly, AppAgent(Zhang et al., 2023) and MobileAgent (Wang et al., 2024a) leverage accessibility APIs to identify GUI components on mobile platforms.
However, as highlighted in UGround (Gou et al., 2024), such metadata can be noisy, incomplete, and computationally expensive to parse at every step. To overcome these limitations, recent research emphasizes visual-only grounding—mapping textual referring expressions or instructions directly to pixel-level coordinates on a screenshot. UGround trains large action models using only screen-level visual inputs. OmniParser (Lu et al., 2024) also demonstrates how vision-only approaches can parse GUIs without HTML or accessibility data. Similarly, OS-Atlas (Wu et al., 2024b) leverages large-scale multi-platform training data to achieve universal GUI grounding that generalizes across web, mobile, and desktop platforms. By unifying data sources and action schemas, OS-Atlas showcases the feasibility of a universal approach to action grounding.
5 GUI Agent Training Methods
This section summarizes different strategies to elicit the ability to solve agentic tasks in a GUI Agent agent. We broadly categorize these strategies into two types: (1) Prompt-based Methods and (2) Training-based Methods. Prompt-based methods do not involve the training of parameters; they elicit the ability to solve agentic tasks by providing detailed instructions or demonstrations within the prompt. Training-based methods, on the other hand, involve optimizing the agent’s parameters to maximize an objective, such as pretraining, fine-tuning, or reinforcement learning.
5.1 Prompt-based Methods
Prompt-based methods enable GUI agents to exhibit learning and adaptation during inference through carefully designed prompts and interaction mechanisms, without modifying model parameters. This learning and adaptation occur as the agent’s state evolves by incorporating context from past actions or stored knowledge.
One key approach is the use of dynamic action generation and accumulation. DynaSaur (Nguyen et al., 2024) enables agents to dynamically create and compose actions by generating and executing Python code via prompting. Given task instructions, the agent outputs code snippets defining new actions or reusing existing ones, effectively learning new skills and improving performance over time. Agent Q (Putta et al., 2024) and OSCAR (Wang and Liu, 2024) incorporate self-reflection and self-critique mechanisms via prompts, enabling agents to iteratively improve decision-making by identifying and rectifying errors. Auto-Intent (Kim et al., 2024) focuses on unsupervised intent discovery and utilization, extracting intents from interaction histories and incorporating them into future prompts. Other techniques include state-space exploration in LASER (Ma et al., 2023), state machine in OSCAR (Wang and Liu, 2024), expert development and multi-agent collaboration in MobileExperts (Zhang et al., 2024b), and app memory in AutoDroid (Wen et al., 2024).
Despite the potential of prompt-based methods, the limited context size of LLMs and the difficulty of designing effective prompts that elicit the desired behavior remain.
5.2 Training-based Methods
5.2.1 Pre-training
Earlier models for GUI tasks relied on assembling smaller encoder-decoder architectures to address visual understanding challenges due to its ability to learn unified representations from diverse visual and textual data, enhance transfer learning capabilities, and integrate multiple modalities deeply. For example, PIX2STRUCT Lee et al. (2023) is pre-trained on a screenshot parsing task, which involves predicting simplified HTML representations from screenshots with visually masked regions. It employs a ViT Dosovitskiy (2020) as the image encoder, T5 Raffel et al. (2020) as the text encoder, and a Transformer-based decoder.
Training of recent GUI agent models often involve the continual pre-training of existing vision large language models on additional large-scale datasets. This step refines the model’s general knowledge and modifies or assembles new neural network modules into the backbone, providing a stronger foundation before fine-tuning on smaller, curated datasets for GUI tasks. VisionLLM Wang et al. (2023) utilizes public datasets to integrate BERT Devlin (2018) and Deformable DETR Zhu et al. (2020) into large language models, focusing on visual question answering tasks centered on grounding and detection. SeeClick Cheng et al. (2024a) is built using continual pre-training on Qwen-VL Bai et al. (2023) with datasets incorporating OCR-based layout annotation to predict click actions. UGround Gou et al. (2024) use continual pre-training on the LLaVA-NEXT Liu et al. (2024a) model without its low-resolution image fusion module on a large dataset and synthetic data to align visual elements with HTML metadata for planning and grounding tasks.
Pre-training is also used to adapt new designs for improved computational efficiency in GUI-related tasks. CogAgent Hong et al. (2023) employs a high-resolution cross-module to process small icons and text, enhancing its efficiency for GUI tasks such as DOM element generation and action prediction. ShowUI Lin et al. (2024) built on Qwen2-VL Wang et al. (2024c) with a visual-token selection module to improve the computational efficiency for interleaved high-resolution grounding.
5.2.2 Fine-tuning
Fine-tuning has emerged as a key strategy to adapt large vision-language models (VLMs) and large language models (LLMs) to the specialized domain of GUI interaction. Unlike zero-shot or prompt-only approaches, fine-tuning can enhance both the model’s grounding in GUI elements and its ability to execute instructions reliably.
Recent work highlights reducing hallucinations and improving grounding. Falcon-UI Shen et al. (2024a) fine-tunes on large-scale instruction-free GUI data and then fine-tunes on Android and Web tasks, achieving high accuracy with fewer parameters. VGA Ziyang et al. (2024), through image-centric fine-tuning, reduces hallucinations by tightly coupling visual inputs with GUI elements, thus improving action reliability. Similarly, UI-Pro Li et al. (2024) identifies a hidden recipe for systematic fine-tuning of VLMs, scaling down model size while maintaining state-of-the-art grounding accuracy.
Other methods leverage fine-tuning to incorporate domain-specific reasoning and functionalities such as functionality-aware fine-tuning for generating human-like interactions Liu et al. (2024d), alignment strategies to handle multilingual, variable-resolution GUI inputs  Nong et al. (2024). Some methods emphasize autonomous adaptation, such as learning to execute arbitrary voice commands through trial-and-error exploration Pan et al. (2023) and learning for cross-platform GUI grounding without structured text Cheng et al. (2024a). Additionally, fine-tuning can specialize models for context-sensitive actions. Techniques proposed by Liu et al. (2023) enable context-aware text input generation, improving coverage in GUI testing scenarios. Taken together, these fine-tuning methods demonstrate how careful parameter adaptation, data scaling and multimodal alignment can collectively advance the reliability, interpretability, and performance of GUI agents.
5.2.3 Reinforcement Learning
Reinforcement learning (RL) was used in the early text-based agent WebGPT to improve information retrieval of the GPT-3 based model Nakano et al. (2021).  Liu et al. (2018) use human demonstrations to constrain the search space for RL, though using workflows as a high-level process for the model to complete without specifying the specific details. An example from  Liu et al. (2018) is for the specific process of forwarding a given email, the workflow would involve clicking forward, typing in the address, and clicking send.  Deng et al. (2023) uses RL based on human demonstrations as the reward signal. While early agents constrained the input and action spaces to only text, recent work has extended to GUI agents.
WebRL framework uses RL to generate new tasks based on previously unsuccessful attempts as a mitigation for sparse rewards Qi et al. (2024). Task success is evaluated by an LLM-based outcome reward model (ORM) and KL-divergence is used to prevent significant shifts in policies during the curriculum. AutoGLM apply online, curriculum learning, in particular to address error recovery during real-world use and to correct for stochasticity not present in simulators Liu et al. (2024c). DigiRL uses a modified advantage-weighted regression (AWR) algorithm for offline learning Peng et al. (2019), but modifies AWR for more stochastic environments by using a simple value function and curriculum learning.
6 Open Problems & Challenges
User Intent Understanding.
GUI Agents still struggle to accurately infer user goals across diverse applications, achieving only 51.1% accuracy on unseen websites Kim et al. (2024). Designing models that generalize effectively across varying tasks is crucial, particularly for handling contextual variations in user interactions Stefanidi et al. (2022) and predicting user behavior in complex interfaces Gao et al. (2024). Future research prospects involve leveraging robust training techniques to enable agents to adapt to new environments with minimal retraining, ultimately providing more seamless and adaptive user experiences.
Security and Privacy.
GUI agents often interact with sensitive data, including passwords, confidential documents, and personal credentials, raising significant privacy and security concerns He et al. (2024a); Zhang et al. (2024a). These risks are amplified when agents rely on cloud-based processing, which involves transmitting sensitive data to remote servers. Unauthorized access or incorrect information could lead to severe consequences Zhang et al. (2024c). Future research could focus on developing privacy-preserving protocols, exploring local processing alternatives, and implementing advanced authentication methods to mitigate these risks and ensure the reliability and safety of GUI agents across diverse environments.
Inference Latency.
The need to handle complex interactions across diverse applications often conflicts with the demand for real-time responsiveness. Optimizing model efficiency without sacrificing accuracy remains a key hurdle, especially when deploying agents in resource-constrained environments. Challenges include reducing computational overhead, integrating hardware acceleration, and balancing trade-offs between speed and resource usage. Addressing these issues will require lightweight model designs and adaptive techniques to ensure timely, seamless interactions in dynamic GUI settings.
7 Conclusion
In this survey, we have thoroughly explored GUI Agents, examining various benchmarks, agent architectures, and training methods. Although considerable strides have been made, problems such as intent understanding, security, latency, and personalization remain critical challenges. We hope this survey will act as a valuable resource for researchers, offering structure and practical guidance in this rapidly growing and exciting field, and inspiring further inquiry into GUI Agents. We are confident that the progress in this area will mark an important milestone, benefiting humankind, significantly enhancing our daily productivity, and transforming the way we interact with computers.
Limitations
We recognize that some studies have explored interactions between LFM-based agents and digital systems through interfaces other than GUIs, such as Command Line Interfaces (CLI) or Application Programming Interfaces (API). However, these approaches are relatively limited in scope compared to GUI-based methods. To maintain a focused scope for our survey, we have chosen not to include them in our discussion.

### v2

@@
 Recently, we have witnessed a paradigm shift from using LFMs purely as conversational chatbots (Touvron et al., 2023; Chiang et al., 2023; Dam et al., 2024) to employing them for performing actions and automating useful tasks (Wang et al., 2024b; Zhao et al., 2023; Yao et al., 2023; Shinn et al., 2023; Shen et al., 2024b; Cheng et al., 2024c).
-In this direction, one approach stands out: leveraging LFMs to interact with digital systems, such as desktops and mobile phones, or software applications such as a web browser, through Graphical User Interfaces (GUIs) in the same way humans do—for example, by controlling the mouse and keyboard to interact with visual elements displayed on a device’s monitor Iong et al.
+In this direction, one approach stands out: leveraging LFMs to interact with digital systems, such as desktops and mobile phones, or software applications such as a web browser, through Graphical User Interfaces (GUIs) in the same way humans do, for example, by controlling the mouse and keyboard to interact with visual elements displayed on a device’s monitor Iong et al.
 (2024); Hong et al.
@@
 This approach holds great potential, as GUIs are ubiquitous across almost all computer devices that humans interact with in their work and daily lives.
-However, deploying LFMs in such environments poses unique challenges, such as dynamic layouts, diverse graphical designs across different platforms, and grounding issues—for instance, fine-grained recognition of elements within a page that are often small, numerous, and scattered Liu et al.
+However, deploying LFMs in such environments poses unique challenges, such as dynamic layouts, diverse graphical designs across different platforms, and grounding issues, for instance, fine-grained recognition of elements within a page that are often small, numerous, and scattered Liu et al.
 (2024b).
@@
 Organization of this Survey.
-We begin our survey by clearly defining the term “GUI Agent,” followed by a traditional RL formalism of GUI Agent tasks in Section 2.
+We begin our survey by clearly defining the term “GUI Agent,” followed by a formal definition of GUI agent tasks in Section 2.
 We then summarize different datasets and environments in Section 3 to provide readers a clearer picture of the kinds of problem settings currently available.
@@
 2 Preliminaries
+This section formally defines the term “GUI agent” and presents a formalization of GUI agent tasks.
 Definition 1 (GUI Agent).
@@
 Problem Formulation.
-GUI Agent involves an agent interacting with an environment in a sequential manner.
-The environment can generally be modeled as a Partially Observable Markov Decision Process (POMDP), defined by a tuple $(\mathcal{U},\mathcal{A},\mathcal{S},\mathcal{O},T)$, where $\mathcal{U}$ is the task space, $\mathcal{A}$ is the action space, $\mathcal{S}$ is the state space, $\mathcal{O}$ is the observation space, and $T:\mathcal{S}\times\mathcal{A}\to\mathcal{P}(\mathcal{S})$ is a state transition function mapping a state-action pair to a probability distribution over subsequent states.
-A GUI Agent is a mapping $\pi:\mathcal{O}\to\mathcal{A}$.
+GUI agent tasks involve an agent interacting with an environment in a sequential manner.
+The environment can generally be modeled as a Partially Observable Markov Decision Process (POMDP) (Sondik, 1971; Hauskrecht, 2000), defined by a tuple $(\mathcal{U},\mathcal{A},\mathcal{S},\mathcal{O},T)$, where $\mathcal{U}$ is the task space, $\mathcal{A}$ is the action space, $\mathcal{S}$ is the state space (not fully observable to the agent), $\mathcal{O}$ is the observation space, and $T:\mathcal{S}\times\mathcal{A}\to\mathcal{P}(\mathcal{S})$ is a state transition function that maps a state-action pair to a probability distribution over subsequent states.
+A GUI agent is a policy $\pi:\Delta^{\mathcal{S}}\to\mathcal{A}$, where $\Delta^{\mathcal{S}}$ denotes the probability simplex over the state states.
+Most commonly, this is implemented using the entire history of past actions and observations.
 Given a task $u\in\mathcal{U}$, the agent proceeds through a sequence of actions to complete the task.
-At each step $t$, based on the available observation $o\in\mathcal{O}$, the agent $\pi$ must predict the next action $a\in\mathcal{A}$.
+At each step $t$, based on the history of past actions and observations, the policy $\pi$ selects the next action $a\in\mathcal{A}$.
 The environment then transitions to a new state $s^{\prime}\in\mathcal{S}$ according to $T$.
-Depending on the design of the environment, the agent may receive a reward $r=R(s,a,s^{\prime})$, where $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a reward function.
+Depending on the environment’s design, the agent may receive a reward $r=R(s,a,s^{\prime})$, where $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a reward function.
 3 Benchmarks
@@
 A GUI environment includes the GUI interface of a mobile phone or a desktop.
-Unlike datasets, environments are dynamic—actions taken within the environment can alter its state, hence, allowing modeling the problem as Markov Decision Processes (MDPs) or Partially Observable MDPs (POMDPs), with defined action, state, and observation spaces, and a state transition function.
+Unlike datasets, environments are dynamic, actions taken within the environment can alter its state, hence, allowing modeling the problem as Markov Decision Processes (MDPs) or Partially Observable MDPs (POMDPs), with defined action, state, and observation spaces, and a state transition function.
 Another critical dimension of the existing benchmarks for GUI agents is the distinction between the open-world and closed-world assumptions.
@@
 In contrast, open-world benchmarks relax this constraint, allowing relevant information required to complete a task to exist outside the benchmark.
+We present a summary of existing GUI agent benchmarks in Table 1.
 3.1 Static Datasets
-3.1.1 Closed-World Datasets
+3.1.1 Closed-World Datasets.
 RUSS dataset introduces real-world instructions mapped to a domain-specific language (DSL) that enables agents to execute web-based tasks with high precision (Xu et al., 2021).
@@
 WebLINX emphasizes multi-turn dialogue for interactive web navigation on real-world sites, enhancing agents’ adaptability and conversational skills (Lù et al., 2024).
-Evaluation on static open-world datasets balances the ease of setting up an evaluation setting with realism since the agents interact with real-world websites.
+Evaluation on static open-world datasets balances the ease of evaluation with realism since the agents interact with real-world websites.
 However, due to the nature of real-world websites, they are often unpredictable and prone to changes, which makes it more challenging to reproduce and compare with prior methods.
@@
 Shifting to enterprise settings, WorkArena evaluates agent performance in complex UI environments, focusing on knowledge work tasks in ServiceNow platform (Drouin et al., 2024).
+WorkArena++ extends this benchmark by introducing more challenging tasks (Boisvert et al., 2024).
 ST-WebAgentBench incorporates safety and trustworthiness metrics, assessing policy adherence and minimizing risky actions, critical for business deployment (Levy et al., 2024).
-Lastly, VideoWebArena introduces long-context video-based tasks, requiring agents to understand instructional videos and integrate them with textual and visual data to complete tasks.
+VideoWebArena introduces long-context video-based tasks, requiring agents to understand instructional videos and integrate them with textual and visual data to complete tasks.
 It emphasizes memory retention and multimodal reasoning (Jang et al., 2024).
+In simulated desktop environments, OSWorld (Xie et al., 2024) provides the first realistic operating system setting for evaluating multimodal GUI agents.
+Spider2-V (Cao et al., 2024) builds on this direction by targeting professional-level data science and engineering tasks.
+BrowserGym (Chezelles et al., 2024) develops a unified ecosystem consisting of seven web agent benchmarks for developing and evaluating web agents.
 Closed-world environments serve as evaluation platforms that mimic the dynamism of real-world environments while offering stability and reproducibility.
@@
 3.3 Evaluation Metrics
-Task Completion Metrics.
+3.3.1 Task Completion Metrics.
 The majority of benchmarks use task completion rate as the primary metric to measure GUI agents’ performance.
@@
 WebVoyager uses GPT-4V to automatically determine success based on the agent’s trajectory, reporting a high agreement rate of 85.3% with human judgments (He et al., 2024b).
-Instead of using a single final-state success metric, WebLINX measures an overall success rate based on aggregated turn-level success rates across tasks (Lù et al., 2024).
-The turn-level success rates are computed depending on the type of actions, e.g., Intersection Over Union (IoU) for click or submit actions, and F1 for say or textinput actions.
-Lastly, there are task-specific metrics to measure success, e.g., using ROUGE-L, F1 for open-ended generation (Liu et al., 2024b; Xu et al., 2024; Wornow et al., 2024), accuracy for multiple choice question tasks (Liu et al., 2024b), Precision and Recall for Standard Operating Procedure (SOP) validation (Wornow et al., 2024), and so on.
-Intermediate Step Metrics.
-While the task completion rate is a straightforward single-numeric metric that simplifies comparing the overall performance of agents, it fails to provide clear insights into their specific behaviors.
+Instead of using a single final-state success metric, WebLINX measures an overall success rate based on aggregated turn-level success metrics across tasks (Lù et al., 2024).
+These turn-level metrics, including Intersection over Union and F1, are computed based on the type of action taken.
+Lastly, there are task-specific metrics to measure success, e.g., using ROUGE-L, F1 for open-ended generation (Liu et al., 2024b; Xu et al., 2024; Wornow et al., 2024), accuracy for multiple choice question tasks (Liu et al., 2024b), Precision and Recall for Standard Operating Procedure validation (Wornow et al., 2024).
+3.3.2 Intermediate Step Metrics.
+While the task completion rate is a single straightforward metric that simplifies evaluation, it fails to provide clear insights into their specific behaviors.
 Although some fine-grained metrics measure step-wise performance, their scope remains limited.
-WebCanvas evaluates step scores using three distinct targets: URL Matching, which verifies whether the agent navigated to the correct webpage; Element Path Matching, which checks if the agent interacted with the appropriate UI element, such as a button or text box; and Element Value Matching, which ensures the agent inputted or extracted the correct values, such as filling a form or reading text.
-WebLINX uses an intent match metric to assess whether the predicted action’s intent aligns with the reference intent.
-Similarly, Mind2Web and MT-Mind2Web evaluate Element Accuracy by measuring the rate at which the agent selects the correct elements.
+WebCanvas (Pan et al., 2024) evaluates step scores using three distinct targets: URL Matching, which verifies whether the agent navigated to the correct webpage; Element Path Matching, which checks if the agent interacted with the appropriate UI element, such as a button or text box; and Element Value Matching, which ensures the agent inputted or extracted the correct values, such as filling a form or reading text.
+WebLINX (Lù et al., 2024) uses an intent match metric to assess whether the predicted action’s intent aligns with the reference intent.
+Similarly, Mind2Web (Deng et al., 2023) and MT-Mind2Web (Deng et al., 2024) evaluate Element Accuracy by measuring the rate at which the agent selects the correct elements.
 These systems also measure the precision, recall, and F1 score for token-level operations, such as clicking or typing, and calculate the Step Success Rate, which reflects the proportion of individual task steps completed correctly.
 While step-wise evaluations provide more fine-grained insight into the agent’s performance, it is often challenging to collect reference labels at the step level while also providing enough flexibility to consider different paths to achieve the original tasks.
-Efficiency, Generalization, Safety and Robustness Metrics.
+3.3.3 Efficiency, Generalization, Safety and Robustness Metrics.
 Lastly, we summarize additional metrics that evaluate various aspects of GUI agents beyond their raw performance.
@@
 4 GUI Agent Architectures
-This section focuses on various architectural designs of a GUI Agent agent, which we categorize into four main types: (1) Perception: designs that enable the GUI Agent agent to perceive and interpret observations from its environment; (2) Reasoning: designs related to the cognitive processes of a GUI Agent agent, such as using an external knowledge base for long-term memory access or a world model of the environment to support other modules like planning; (3) Planning: designs related to decomposing a task into subtasks and creating a plan for their execution; and (4) Acting: mechanisms that allow the GUI Agent agent to interact with the environment, including representing actions in natural language using specific templates, JSON, or programming languages as action representations.
+This section focuses on various architectural designs of a GUI agent, which we categorize into four main types: (1) Perception: designs that enable the GUI agent to perceive and interpret observations from its environment; (2) Reasoning: designs related to the cognitive processes of a GUI agent, such as using an external knowledge base for long-term memory access or a world model of the environment to support other modules like planning; (3) Planning: designs related to decomposing a task into subtasks and creating a plan for their execution; and (4) Acting: mechanisms that allow the GUI agent to interact with the environment, including representing actions in natural language using specific templates, JSON, or programming languages as action representations.
+We present a taxonomy of GUI agent architectures in Figure 1.
 4.1 Perception
-Unlike API-based agents that process structured, program-readable data, GUI agents must perceive and understand the on-screen environment that is designed for human consumption.
+Unlike API-based agents that process structured, program-readable data Xu et al.
+(2025b), GUI agents must perceive and understand the on-screen environment that is designed for human consumption.
 This requires carefully chosen interfaces that allow agents to discover the location, identity, and properties of the interactive elements.
@@
 4.1.3 Screen-visual-based Interfaces
-With advances in computer vision and multimodal LLM, agents can utilize screen-visual information, like screenshots, to perceive on-screen environment.
+With advances in computer vision and multimodal LLMs, agents can utilize screen-visual information, such as screenshots, to perceive the on-screen environment.
 OmniParser (Lu et al., 2024) utilizes an existing multimodal LLM (e.g., GPT-4V) to parse a screenshot into a structured representation of the UI elements.
+TAG (Xu et al., 2025a) leverages the inherent attention patterns in pretrained MLLMs to improve GUI grounding without the need for additional fine-tuning.
+Cradle (Tan et al., 2024), instead of relying on a single screenshot, processes a video recording (i.e., a sequence of screenshots) to enable more general-purpose computer control.
 However, screen-visual-based perception introduces privacy concerns since entire screenshots may contain sensitive information.
@@
 Despite these challenges, such interfaces are crucial for agents operating in environments where high-quality accessibility interfaces and DOM information are unavailable, or environments where dynamic or visual information is crucial, like image or video editing software.
+A key advantage of this approach is that it requires no application instrumentation, enabling direct deployment across a wide range of applications.
 4.1.4 Hybrid Interfaces
-To achieve robust and flexible performance across diverse environments, many GUI agents employ a hybrid approach.
+To achieve robust and flexible performance across diverse environments, many GUI agents employ a hybrid approach (Gou et al., 2024; Wu et al., 2024b; Kil et al., 2024).
 These systems combine accessibility APIs, DOM data, and screen-visual information to form a more comprehensive understanding of the interface.
-Leading methods in GUI agent tasks, such as OS-Atlas(Wu et al., 2024b) and UGround (Gou et al., 2024), demonstrates that hybrid interfaces that combine visual and textual inputs can enhance performance.
-Hybrid interfaces based approaches also facilitate error recovery—when accessibility or DOM data are incomplete or misleading, the agent can fall back on screen parsing, and vice versa.
+Leading methods in GUI agent tasks, such as OS-Atlas (Wu et al., 2024b) and UGround (Gou et al., 2024), demonstrate that hybrid interfaces that combine visual and textual inputs can enhance performance.
+Such approaches also facilitate error recovery, when accessibility or DOM data are incomplete or misleading, the agent can fall back on screen parsing, and vice versa.
 4.2 Reasoning
@@
 (2024).
-Traditional planning methods, such as symbolic approaches and reinforcement learning, have significant limitations: symbolic methods require extensive human expertise to define rigid system rules and lack error tolerance Belta et al.
+Traditional planning methods, such as symbolic approaches (Kautz and Selman, 1992) and reinforcement learning (Sutton and Barto, 1998), have significant limitations: symbolic methods require extensive human expertise to define rigid system rules and lack error tolerance Belta et al.
 (2007); Pallagani et al.
@@
 (2000).
-WebDreamer (Gu et al., 2024) uses LLMs to simulate the outcomes of the actions of each agent and then evaluate the result to determine the optimal plan at each step.
+WebDreamer (Gu et al., 2024) uses LLMs to simulate the outcomes of the actions of each agent and then evaluates the result to determine the optimal plan at each step.
 MobA (Zhu et al., 2024) devises a two-level architecture to power the mobile phone management, with a high level for understanding user commands, tracking history memories and planning tasks, and a low level to act the planned module.
@@
 The aggregation module subsequently consolidates solutions from these analogous problems, enhancing the problem-solving process for the original input.
-WebShop, Mind2Web, and WebArena Zhou et al.
+Benchmarks like WebShop, Mind2Web, and WebArena Zhou et al.
 (2023c); Deng et al.
-(2023) allow agents to interact with webs to plan for web browsing for search.
+(2023) enable agents to interact with web environments to plan and execute browsing actions for information-seeking tasks.
 WMA (Chae et al., 2024) utilizes world models to address the mistakes made by LLMs for long-horizon tasks.
@@
 Unlike purely text-based or API-driven agents, GUI agents must articulate their actions at a finer granularity—often down to pixel-level coordinates—while also handling higher-level semantic actions such as typing text, scrolling, or clicking on specific elements.
-Several directions of approaches have emerged:
+Several approaches have emerged:
 Those utilizing textual interfaces may only rely on text-based metadata (HTML, accessibility trees) to identify UI elements.
@@
 We broadly categorize these strategies into two types: (1) Prompt-based Methods and (2) Training-based Methods.
-Prompt-based methods do not involve the training of parameters; they elicit the ability to solve agentic tasks by providing detailed instructions or demonstrations within the prompt.
+Prompt-based methods do not involve the training of parameters; they elicit the ability to solve agentic tasks by providing detailed instructions within the prompt.
 Training-based methods, on the other hand, involve optimizing the agent’s parameters to maximize an objective, such as pretraining, fine-tuning, or reinforcement learning.
+We present a taxonomy of GUI agent training methods in Figure 2.
 5.1 Prompt-based Methods
@@
 This learning and adaptation occur as the agent’s state evolves by incorporating context from past actions or stored knowledge.
-One key approach is the use of dynamic action generation and accumulation.
-DynaSaur (Nguyen et al., 2024) enables agents to dynamically create and compose actions by generating and executing Python code via prompting.
-Given task instructions, the agent outputs code snippets defining new actions or reusing existing ones, effectively learning new skills and improving performance over time.
 Agent Q (Putta et al., 2024) and OSCAR (Wang and Liu, 2024) incorporate self-reflection and self-critique mechanisms via prompts, enabling agents to iteratively improve decision-making by identifying and rectifying errors.
@@
 UGround Gou et al.
-(2024) use continual pre-training on the LLaVA-NEXT Liu et al.
+(2024) uses continual pre-training on the LLaVA-NEXT Liu et al.
 (2024a) model without its low-resolution image fusion module on a large dataset and synthetic data to align visual elements with HTML metadata for planning and grounding tasks.
@@
 ShowUI Lin et al.
-(2024) built on Qwen2-VL Wang et al.
+(2024) builds on Qwen2-VL Wang et al.
 (2024c) with a visual-token selection module to improve the computational efficiency for interleaved high-resolution grounding.
@@
 Similarly, UI-Pro Li et al.
-(2024) identifies a hidden recipe for systematic fine-tuning of VLMs, scaling down model size while maintaining state-of-the-art grounding accuracy.
-Other methods leverage fine-tuning to incorporate domain-specific reasoning and functionalities such as functionality-aware fine-tuning for generating human-like interactions Liu et al.
-(2024d), alignment strategies to handle multilingual, variable-resolution GUI inputs  Nong et al.
+(2024) identifies a recipe for fine-tuning of VLMs, reducing model size while maintaining state-of-the-art grounding accuracy.
+Other methods leverage fine-tuning to incorporate domain-specific reasoning and functionalities, such as functionality-aware fine-tuning for generating human-like interactions Liu et al.
+(2024d) and alignment strategies to handle multilingual, variable-resolution GUI inputs  Nong et al.
 (2024).
@@
 5.2.3 Reinforcement Learning
-Reinforcement learning (RL) was used in the early text-based agent WebGPT to improve information retrieval of the GPT-3 based model Nakano et al.
+Reinforcement learning was used in the early text-based agent WebGPT to improve information retrieval of the GPT-3 based model Nakano et al.
 (2021).
 Liu et al.
-(2018) use human demonstrations to constrain the search space for RL, though using workflows as a high-level process for the model to complete without specifying the specific details.
+(2018) use human demonstrations to constrain the search space for RL, through using workflows as a high-level process for the model to complete without specifying the specific details.
 An example from  Liu et al.
@@
 Deng et al.
-(2023) uses RL based on human demonstrations as the reward signal.
+(2023) use RL based on human demonstrations as the reward signal.
 While early agents constrained the input and action spaces to only text, recent work has extended to GUI agents.
@@
 (2024).
-Task success is evaluated by an LLM-based outcome reward model (ORM) and KL-divergence is used to prevent significant shifts in policies during the curriculum.
-AutoGLM apply online, curriculum learning, in particular to address error recovery during real-world use and to correct for stochasticity not present in simulators Liu et al.
+Task success is evaluated by an LLM-based outcome reward model (ORM) and KL-divergence is used to prevent significant shifts in policies during curriculum learning.
+AutoGLM applies online, curriculum learning, in particular to address error recovery during real-world use and to correct for stochasticity not present in simulators Liu et al.
 (2024c).
@@
 (2024).
-Future research prospects involve leveraging robust training techniques to enable agents to adapt to new environments with minimal retraining, ultimately providing more seamless and adaptive user experiences.
+A prospective future research direction is to leverage robust training techniques to enable agents to adapt to new environments with minimal retraining, ultimately providing more seamless and adaptive user experiences.
+Other promising directions could include training on diverse user interaction datasets and incorporating context-aware learning techniques that utilize historical user actions to better predict intent.
 Security and Privacy.
-GUI agents often interact with sensitive data, including passwords, confidential documents, and personal credentials, raising significant privacy and security concerns He et al.
+GUI agents frequently interact with sensitive data such as passwords, confidential documents, and personal credentials, raising serious privacy and security concerns He et al.
 (2024a); Zhang et al.
 (2024a).
-These risks are amplified when agents rely on cloud-based processing, which involves transmitting sensitive data to remote servers.
-Unauthorized access or incorrect information could lead to severe consequences Zhang et al.
+These risks are further amplified when agents rely on cloud-based processing, which involves transmitting sensitive information to remote servers.
+Unauthorized access or incorrect actions could result in severe consequences Zhang et al.
 (2024c).
-Future research could focus on developing privacy-preserving protocols, exploring local processing alternatives, and implementing advanced authentication methods to mitigate these risks and ensure the reliability and safety of GUI agents across diverse environments.
+Future research could focus on developing privacy-preserving protocols, such as homomorphic encryption or differential privacy, to ensure data remains secure during both inference and storage.
+Additional directions may include exploring local processing alternatives and implementing advanced authentication mechanisms to enhance the reliability and safety of GUI agents across diverse environments.
 Inference Latency.
-The need to handle complex interactions across diverse applications often conflicts with the demand for real-time responsiveness.
-Optimizing model efficiency without sacrificing accuracy remains a key hurdle, especially when deploying agents in resource-constrained environments.
-Challenges include reducing computational overhead, integrating hardware acceleration, and balancing trade-offs between speed and resource usage.
-Addressing these issues will require lightweight model designs and adaptive techniques to ensure timely, seamless interactions in dynamic GUI settings.
+The need to manage complex interactions across diverse applications often conflicts with the requirement for real-time responsiveness.
+Optimizing model efficiency without compromising accuracy remains a key challenge, particularly when deploying agents in resource-constrained environments.
+Key issues include minimizing computational overhead, leveraging hardware acceleration, and balancing trade-offs between speed and resource usage.
+Addressing these challenges calls for lightweight model architectures and adaptive techniques that enable timely, seamless interactions in dynamic GUI settings.
+Future research could investigate hardware-aware optimization methods, such as quantization and pruning, or efficient decoding strategies like predictive sampling and multi-token prediction, which can significantly reduce latency while preserving system accuracy.
+Personalization.
+is a pivotal aspect in the development of GUI agents, aiming to tailor interactions to individual user preferences and behaviors, thereby enhancing satisfaction and efficiency.
+Recent work Berkovitch et al.
+(2024) introduced a method for identifying user goals from UI trajectories, enabling agents to infer intentions and proactively assist users based on their interactions with the interface.
+Future research could explore more sophisticated models that incorporate user feedback to refine personalization strategies, while ensuring trust and compliance with data protection regulations.
+Additional directions include implementing explicit feedback mechanisms (e.g., thumbs-up/thumbs-down ratings) and developing robust user profiling techniques that integrate behavioral and contextual data to enable more meaningful and adaptive personalization.
 7 Conclusion
@@
 Although considerable strides have been made, problems such as intent understanding, security, latency, and personalization remain critical challenges.
-We hope this survey will act as a valuable resource for researchers, offering structure and practical guidance in this rapidly growing and exciting field, and inspiring further inquiry into GUI Agents.
-We are confident that the progress in this area will mark an important milestone, benefiting humankind, significantly enhancing our daily productivity, and transforming the way we interact with computers.
+We hope that this survey is a valuable resource for researchers, offering structure and practical guidance in this rapidly growing and exciting field, and inspiring more work on GUI Agents.
+The progress in this area has already benefited mankind, enhancing our daily productivity and transforming the way we interact with computers.
 Limitations
-We recognize that some studies have explored interactions between LFM-based agents and digital systems through interfaces other than GUIs, such as Command Line Interfaces (CLI) or Application Programming Interfaces (API).
-However, these approaches are relatively limited in scope compared to GUI-based methods.
+We recognize that some studies have explored interactions between LFM-based agents and digital systems through interfaces other than GUIs, such as Command Line Interfaces (CLI) (Nguyen et al., 2024) or Application Programming Interfaces (API).
+(Song et al., 2025) However, these approaches are relatively limited in scope compared to GUI-based methods.
 To maintain a focused scope for our survey, we have chosen not to include them in our discussion.